An electron in the 4th principal energy level will have energy __________ an electron in the 1 st principal energy level.

higher than

lower than

equal to
Answer #1

The energy of an electron in n-th energy level is , En = -2π2mZ2e4/n2h2

where, m = mass of electron

Z = atomic no of atom, e = charge of an electron, n= principle energy level, h = planck's constant

For a particular atom Z is constant, the other parameter m, e, h are constant

So it can be written as En ∝ - 1/n2

So with increasing no of n the numerical value of En decreases. As there is a negative sign before the the whole expression so overall the value of E4 will be higher than E1

E4 = - E1/42 = -E1/16 so E4 > E1.

Electron will release ( E4-E1) amount of energy when it jumps from E4 level to E1 level

So electron in the 4th principal energy level will have energy higher than an electron in the 1 st principal energy level.